OpenStudy (anonymous):

the region bounded by y=5 and y=x+4/x is rotated about the line x=-1, find the volume the resulting solid by any method

OpenStudy (sidsiddhartha):

first find the intersection of the curves

OpenStudy (sidsiddhartha):

5 = x + 4/x so solve it x^2 - 5x + 4 = 0 (x - 1) (x - 4) = 0 x = 1, 4 now u can solve it using shell method

OpenStudy (sidsiddhartha):

the graph will look like this http://www.wolframalpha.com/input/?i=y%3Dx%2B%284%2Fx%29+and+y%3D5

OpenStudy (sidsiddhartha):

now think the solid as a set of cylindrical shells one inside other now what it the hieght of each shell?

OpenStudy (sidsiddhartha):

it is simply distance between y = 5 and y = x + 4/x

OpenStudy (anonymous):

yeah i know i need to find intersections. the problem that i have is i cant find the correct p(x) and h(x). Could you help me out? Thanks so much.

OpenStudy (sidsiddhartha):

so H=5-x-4/x and radius R=(x+1) now volume \[\int\limits_{1}^{4}2 \pi RH=volume\] \[volume=\int\limits_{1}^{4}2*\pi *(5 - x - 4/x)(x+1)dx\]

OpenStudy (anonymous):

oh i see! i got the wrong boundaries. thanks so much! that really helped!

OpenStudy (sidsiddhartha):

:)